OpenStudy (anonymous):

find the exact value of the given expression Cos(60*+315*)

OpenStudy (anonymous):

it should be the same value as cos(15)

OpenStudy (anonymous):

cos (60+315) = cos(375)= cos(360+15)

OpenStudy (anonymous):

cos (15) = cos(45-30)

OpenStudy (anonymous):

use the formula cos (a-b)= cos a cos b - sin a sin b

OpenStudy (anonymous):

therefore, cos (15) = cos 45 cos 30 - sin 45 sin 30

OpenStudy (anonymous):